OpenStudy (anonymous):

solve the proportion using cross products. Round to the nearest hundredth if necessary. 22 m=12 m 55 h = h hours

OpenStudy (anonymous):

22 mh = 12* 55 mh h = 12* 55 / 22 = 30 hours